OpenStudy (anonymous):

v-w=-5 v+2w=4 Use linear combinations to solve the system of linear equations.

OpenStudy (anonymous):

What does the v-w=-5 stand for

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

if i dont know then i cant solve it, i think

OpenStudy (anonymous):

What does it say in the book

OpenStudy (anonymous):

solve by substitution method for systems of equations: v-w=-5 v+2w=4 v=-5+w (-5+w)+2w=4 -5+3w=4 3w=9 w=3 v-(3)=-5 v=-2 w=3 v=-2